### v1

# Comparative Analysis of QSVM and VQC Models for B-cell Epitope Prediction in Vaccine Design Using Quantum Machine Learning

Abstract
Support Vector Machine (SVM) is a classical and widely applied supervised machine learning algorithm for binary classification. Utilizing the kernel trick, it enables the embedding of data into higher-dimensional feature spaces to effectively address non-linear classification tasks.
This study primarily investigates the classical Support Vector Machine (SVM) and introduces an enhanced quantum version—Quantum Support Vector Machine (QSVM)—constructed using quantum circuits, to compare their performance on classification tasks. In QSVM, the high-dimensional computations typically performed by kernel methods are offloaded to quantum circuits, enabling data to be mapped into higher-dimensional feature spaces and potentially accelerating computation, with the aim of improving overall performance.
In addition, this study incorporates a comparison with the Variational Quantum Classifier (VQC), a purely quantum model that entirely replaces classical computation with quantum circuits. VQC is designed to harness the advantages of quantum computing in handling high-dimensional feature spaces, making it a promising approach for classification and prediction tasks.
In recent years, the outbreak of COVID-19 has underscored the importance of vaccine development. One critical step in this process is the prediction of B-cell epitopes. In this study, a B-cell epitope dataset from the Immune Epitope Database (IEDB) is used for model training and prediction. The dataset contains 10 features that describe each B-cell epitope.
To facilitate the analysis of model performance across different dimensions, Principal Component Analysis (PCA) is applied for dimensionality reduction. This allows us to examine how performance varies from lower- to higher-dimensional feature spaces. Additionally, the effect of sample size on prediction outcomes is investigated, treating it as a variable in the analysis.
QSVM and VQC were found to perform well under different conditions, with VQC showing stronger performance in larger datasets and QSVM being effective in low-sample regimes. This work provides new insights into the practical deployment of quantum machine learning for biomedical sequence classification tasks, especially in vaccine design scenarios.
1 Introduction
The original motivation behind quantum computing was to overcome the inherent limitations of classical computers. When simulating quantum systems on classical machines, the exponential growth of the Hilbert space leads to intractable computational costs. In 1982, Richard Feynman proposed using quantum systems to simulate quantum phenomena, laying the foundation for the field.
Quantum machine learning (QML), which integrates quantum computing with machine learning frameworks, has emerged as a promising field. Two notable quantum classifiers are the quantum support vector machine (QSVM) and the variational quantum classifier (VQC), both of which have shown potential in solving classification problems with high-dimensional data.
Recent studies have demonstrated the application of QSVM and VQC in biomedical domains. Zhuang et al. [25] successfully used a QSVM to classify non-hemolytic peptides, outperforming classical SVM models on biological datasets. Similarly, Li et al. [26] presented a comprehensive review of QML applications in drug discovery and immunoinformatics, reinforcing the role of quantum-enhanced models in medical research.
In this work, we investigate the use of QSVM and VQC for predicting linear B-cell epitopes — a crucial component of vaccine design. Using benchmark datasets from IEDB, we apply PCA-based preprocessing and conduct a detailed comparison of model performance under varying input dimensionality and sample sizes. Our results provide insight into the practical feasibility of QML approaches in computational immunology.
2 Support Vector Machine
2.1 Overview
Support Vector Machines (SVMs) are supervised learning models rooted in statistical learning theory [14]. Their core strength lies in transforming input data into higher-dimensional feature spaces, where a linear separator (hyperplane) can be more easily identified for non-linearly separable data. The optimal decision boundary is determined by support vectors—data points closest to the hyperplane—while the objective is to maximize the margin between classes.
2.2 Kernel Trick and Nonlinear Classification
In real-world scenarios, data is rarely linearly separable. To address this, SVMs utilize the kernel trick, which implicitly maps input data to a high-dimensional feature space via a kernel function $K(x_{i},x_{j})=\phi(x_{i})\cdot\phi(x_{j})$, where $\phi$ is a feature map into a Hilbert space $H$. This allows SVMs to construct nonlinear decision boundaries while maintaining computational efficiency.
Let $T=\{(x_{1},y_{1}),(x_{2},y_{2}),\dots,(x_{t},y_{t})\}$, where $x_{i}\in\mathbb{R}^{n}$, $y_{i}\in\{-1,+1\}$. The learning objective is to identify a hyperplane in the transformed feature space that best separates the classes, formalized as a convex quadratic programming problem.
2.3 Relevance to Quantum Models
The QSVM used in this work inherits the core ideas of classical SVMs—particularly margin maximization and the kernel method—but implements them within a quantum computing framework. This enables efficient handling of large-dimensional feature spaces using quantum state representations and inner-product evaluations.
3 Background and Data Preprocessing
To gain an intuitive understanding of the dataset structure after dimensionality reduction, we applied PCA to the original feature space and visualized the first two and three principal components. As shown in Figures 1 and 2, the projected data displays partial separability between positive and negative classes, suggesting that low-dimensional representations may still preserve useful discriminative features.
Figure 1: 2D PCA projection of the input features.
Figure 2: 3D PCA projection of the input features.
3.1 B-cell Epitope Prediction Background
B cells play a central role in the adaptive immune system by recognizing antigenic determinants (epitopes) on protein surfaces and producing antibodies specific to those regions. Accurate identification of linear B-cell epitopes is essential for rational vaccine design, as it enables the selection of high-immunogenicity protein segments for epitope-based vaccines.
To address the need for efficient and scalable epitope prediction, especially in the context of emergent viral threats such as SARS-CoV-2, immunoinformatics approaches have become increasingly prevalent. Databases such as the Immune Epitope Database (IEDB) and UniProt provide well-annotated repositories of peptide sequences and their associated immune activities. In our study, we focus on linear B-cell epitope data extracted from IEDB, specifically those associated with IgG-class antibodies, which represent the majority of curated entries.
3.2 Dataset Description and Preprocessing
The dataset comprises amino acid sequences annotated with binary activity labels: ”positive” (inducing an immune response) and ”negative” (non-inducing). To maintain data consistency and avoid measurement bias, records with conflicting or ambiguous quantitative annotations were excluded.
Each epitope candidate is described using ten physicochemical descriptors that encode amino acid properties relevant to immunogenicity. Prior to model training, the dataset is standardized and undergoes dimensionality reduction using Principal Component Analysis (PCA). This step ensures that models are trained under various levels of feature complexity, allowing us to explore performance across different dimensional regimes.
3.3 Label Distribution and Evaluation Metrics
The final dataset is balanced to the extent possible, and divided into training and testing subsets. We adopt standard metrics to evaluate classifier performance, including Accuracy (ACC), Area Under the ROC Curve (AUC), and the Matthews Correlation Coefficient (MCC). These metrics offer complementary insights: ACC provides overall correctness, AUC measures separability between classes, and MCC accounts for class imbalance.
4 Quantum Support Vector Machine, QSVM
Figure 3: Configuration diagram of the Quantum Support Vector Machine (QSVM).
4.1 Kernel-based Classification Using Quantum Circuits
In classical SVMs, nonlinear classification is enabled through the use of kernel functions, which compute inner products in a high-dimensional feature space:
$K(x_{i},x_{j})=\langle\phi(x_{i}),\phi(x_{j})\rangle$
where $\phi$ is a nonlinear feature map and $K$ represents the kernel matrix.
The Quantum Support Vector Machine (QSVM) extends this concept by encoding data into quantum states and using a quantum circuit to compute the kernel matrix. A typical QSVM uses the fidelity (overlap) between quantum states as the kernel value:
$K(x_{i},x_{j})=|\langle\phi(x_{i})|\phi(x_{j})\rangle|^{2}$
The quantum feature maps $\phi(x)$ are implemented by parametrized quantum circuits. The resulting kernel matrix is estimated by repeatedly executing the circuit and measuring in the computational basis. The number of times the all-zero outcome is observed determines the kernel estimate. This quantum kernel is then passed to a classical SVM solver for final classification, forming a hybrid quantum-classical architecture.
4.2 Data Encoding and Feature Mapping
For quantum computation, classical data must first be embedded into quantum states. This is achieved using a feature map $\phi(x)$, implemented via a quantum circuit. In this study, we adopt angle encoding, where each input feature is mapped to a rotational angle on a qubit using single-qubit rotation gates. Given $n$ features, the input vector $x\in\mathbb{R}^{n}$ is encoded as:
$|x\rangle=U_{\phi}(x)|0\rangle^{\otimes n}$
where $U_{\phi}(x)$ is a unitary operation encoding the data.
The expressiveness of the feature map significantly impacts the performance of the QSVM. Deeper or entangling circuits may yield richer feature spaces but also incur higher quantum hardware demands.
5 Variational Quantum Classifier, VQC
Figure 4: Configuration diagram of the Variational Quantum Classifier (VQC).
5.1 Concept and Architecture
The Variational Quantum Classifier (VQC) is a hybrid quantum-classical model designed for classification tasks. It consists of two main components: a quantum circuit parameterized by rotation and entangling gates (called a variational form), and a classical optimizer used to minimize a cost function based on prediction error.
During training, the input data is first encoded into quantum states via a feature map. These encoded states are processed by the variational circuit, and measurements are performed to extract class probabilities. The parameters of the variational circuit are updated iteratively using a classical optimization algorithm (e.g., COBYLA, SPSA) to minimize the empirical risk.
5.2 Implementation Details
In our implementation, we adopt angle encoding as the input embedding scheme. The variational form comprises parameterized single-qubit $R_{y}$ rotations and entangling $CX$ gates across $n$ qubits. The expectation values of Pauli-Z measurements on each qubit are aggregated and post-processed using a thresholding rule to determine the predicted class label.
This approach allows for end-to-end training of quantum circuits in a supervised learning setting. While VQC models can be expressive, they are also susceptible to issues such as barren plateaus, where the gradients vanish and training stagnates—this remains an active area of research.
6 Performance Evaluation
6.1 QSVM and VQC Experimental Results
B-cell epitope prediction datasets from IEDB and UniProt were used in this study. Each sample consists of 10 features, and the dataset contains over 14,000 entries. Of these, more than 10,000 are labeled as positive, and approximately 4,000 are labeled as negative.
Before applying support vector machine classification to the dataset, we performed dimensionality reduction using PCA, followed by normalization using the StandardScaler from the scikit-learn library. PCA was used to reduce the original 10-dimensional data to a range of 2 to 10 dimensions for analysis. We also investigated how varying the number of training samples—from small to large—affects the model’s performance. The QSVM was implemented using Qiskit on IBM’s quantum computing platform.
6.2 QSVM
Table 2 shows the results of the QSVM. The highest accuracy, 70%, was achieved when using 20 samples with 2-dimensional input features.
Dimensions Sample size 20 100 Dimensions Sample size 20 100
2 0.7 0.69 7 0.55 0.69
3 0.55 0.69 8 0.55 0.69
4 0.45 0.69 9 0.55 0.69
5 0.69 0.69 10 0.55 0.69
6 0.6 0.69
Table 1: QSVM Accuracy
Table 3 presents the AUC values of the QSVM for sample sizes of 20 and 100, with input dimensions ranging from 2 to 10.
Dimensions Sample size 20 100
2 0.7071 0.5
3 0.5404 0.4871
4 0.4191 0.5996
5 0.5 0.5
6 0.5 0.5
7 0.5 0.5
8 0.5 0.5
9 0.5 0.5
10 0.5 0.5
Table 2: QSVM AUC value
According to Table 3, the highest AUC value of 0.7071 was observed when the input dimension was set to 2 and the sample size was 20.
Table 4 presents the MCC values of the QSVM for sample sizes of 20 and 100, with input dimensions ranging from 2 to 10.
Dimensions Sample size 20 100
2 0.4141 0.0
3 0.0820 -0.0545
4 -0.0201 0.0475
5 0.0 0.0
6 0.0 0.0473
7 0.0 0.0
8 0.0 0.0
9 0.0 0.0
10 0.0 0.0
Table 3: QSVM MCC value
According to Table 4, the highest MCC value of 0.4141 was achieved with an input dimension of 2 and a sample size of 20.
6.3 The Variational Quantum Classifier
For the Variational Quantum Classifier (VQC), we similarly explored the effect of varying input dimensions from 2 to 10 and different sample sizes. Since the VQC relies on a variationally optimized circuit, we conducted experiments using 10, 100, and 150 training epochs to analyze its performance under different optimization depths.
Table 5 presents the accuracy of the VQC with 10 training epochs, where the sample size ranges from 1,000 to 4,000 and the input dimension varies from 2 to 10.
Dimensions Sample size 1000 2000 3000 4000
2 0.7300 0.7230 0.7293 0.7260
3 0.7300 0.7230 0.7293 0.7260
4 0.7350 0.7245 0.7333 0.7280
5 0.7310 0.7220 0.7303 0.7258
6 0.7330 0.7250 0.7300 0.7253
7 0.7280 0.7245 0.7320 0.7240
8 0.7280 0.7240 0.7290 0.7260
9 0.7300 0.7240 0.7293 0.7262
10 0.7300 0.7235 0.7320 0.7275
Table 4: VQC 10 epoch Accuracy
As shown in Table 5, the accuracy generally hovers around 73%, with the highest accuracy achieved when the sample size is 1,000 and the input dimension is 4.
Table 6 presents the AUC values for 10 training epochs, with sample sizes ranging from 1,000 to 4,000 and input dimensions from 2 to 10.
Dimensions Sample size 1000 2000 3000 4000
2 0.5417 0.5090 0.5651 0.5587
3 0.5111 0.5519 0.5765 0.5736
4 0.5595 0.5855 0.6090 0.6210
5 0.6069 0.5813 0.6483 0.6245
6 0.6251 0.6211 0.6593 0.6409
7 0.5915 0.5880 0.6545 0.6533
8 0.5412 0.6288 0.6602 0.6587
9 0.5687 0.6122 0.6555 0.6467
10 0.5909 0.6289 0.6730 0.6750
Table 5: VQC 10 epoch AUC value
As shown in Table 6, the best AUC value of 0.6750 was obtained when the sample size was 4,000 and the input dimension was 10.
Table 7 presents the MCC values for 10 training epochs, with sample sizes ranging from 1,000 to 4,000 and input dimensions from 2 to 10.
Dimensions Sample size 1000 2000 3000 4000
2 0.0 0.0 0.0 0.0
3 0.0 0.0 0.0 0.0
4 0.0110 0.0644 0.0987 0.0729
5 0.0520 0.0069 0.0597 0.0100
6 0.0851 0.0686 0.0399 0.0011
7 -0.0029 0.0626 0.0943 -0.0079
8 -0.0272 0.0511 -0.0111 0.0
9 0.0 0.0511 0.0327 0.0241
10 0.0 0.0337 0.0792 0.0585
Table 6: VQC 10 epoch MCC value
In Table 7, the highest MCC value of 0.0987 was achieved when the sample size was 3,000 and the input dimension was 4.
Table 8 presents the accuracy results for 100 training epochs, with sample sizes ranging from 1,000 to 4,000 and input dimensions from 2 to 10.
Dimensions Sample size 1000 2000 3000 4000
2 0.7290 0.7245 0.7293 0.7260
3 0.7290 0.7320 0.7293 0.7260
4 0.7330 0.7320 0.7337 0.7272
5 0.7320 0.7320 0.7327 0.7260
6 0.7370 0.7240 0.7300 0.7258
7 0.7280 0.7230 0.7310 0.7253
8 0.7170 0.7230 0.7313 0.7270
9 0.7330 0.7230 0.7310 0.7268
10 0.7290 0.7230 0.7357 0.7272
Table 7: VQC 100 epoch Accuracy
The accuracy values in Table 8 are similar to those observed at 10 epochs, remaining around 73%. In this case, the highest accuracy was achieved with a sample size of 3,000 and an input dimension of 10.
Table 9 presents the AUC values for 100 training epochs, with sample sizes ranging from 1,000 to 4,000 and input dimensions from 2 to 10.
Dimensions Sample size 1000 2000 3000 4000
2 0.5726 0.5441 0.5771 0.5713
3 0.5731 0.5599 0.5953 0.5804
4 0.5914 0.5996 0.6245 0.6334
5 0.6198 0.6053 0.6641 0.6668
6 0.6525 0.6460 0.6730 0.6645
7 0.6230 0.6612 0.6714 0.6673
8 0.6007 0.6540 0.6814 0.6725
9 0.6225 0.6271 0.6791 0.6774
10 0.6619 0.6395 0.6982 0.7000
Table 8: VQC 100 epoch AUC value
In Table 9, the AUC value reaches 0.7, which occurs at an input dimension of 10 with a sample size of 4,000—the largest sample size and the highest dimensional setting in the experiment.
Table 10 presents the MCC values for 100 training epochs, with sample sizes ranging from 1,000 to 4,000 and input dimensions from 2 to 10.
Dimensions Sample size 1000 2000 3000 4000
2 0.0374 0.0626 0.0 0.0
3 0.0780 0.0 0.0 0.0
4 0.0851 0.0475 0.0987 0.1008
5 0.0718 0.0 0.0973 0.0
6 0.1298 0.0473 0.0548 0.0181
7 0.0532 0.0 0.0802 0.0142
8 0.0376 0.0 0.0716 0.0483
9 0.0846 0.0 0.0702 0.0446
10 0.0374 0.0 0.0792 0.1224
Table 9: VQC 100 epoch MCC value
As shown in Table 10, the MCC value exceeds 0.1 when the sample size is 1,000 and the input dimension is 6, representing a relatively good performance.
Table 11 presents the accuracy results for 150 training epochs, with sample sizes ranging from 1,000 to 4,000 and input dimensions from 2 to 10.
Dimensions Sample size 1000 2000 3000 4000
2 0.7290 0.7245 0.7293 0.7260
3 0.7290 0.7230 0.7293 0.7260
4 0.7310 0.7320 0.7347 0.7272
5 0.7290 0.7245 0.7327 0.7260
6 0.7330 0.7240 0.7280 0.7260
7 0.7310 0.7240 0.7310 0.7255
8 0.7270 0.7240 0.7320 0.7270
9 0.7310 0.7230 0.7343 0.7270
10 0.7250 0.7230 0.7383 0.7272
Table 10: VQC 150 epoch Accuracy
Compared to the previous two epoch settings (Tables 5 and 8), Table 11 shows a slight increase in accuracy, reaching nearly 74
Table 12 presents the AUC values for 150 training epochs, with sample sizes ranging from 1,000 to 4,000 and input dimensions from 2 to 10.
Dimensions Sample size 1000 2000 3000 4000
2 0.5726 0.5441 0.5771 0.5713
3 0.5721 0.5599 0.5953 0.5804
4 0.5809 0.5996 0.6315 0.6334
5 0.6182 0.6084 0.6624 0.6634
6 0.6519 0.6460 0.6762 0.6603
7 0.6378 0.6722 0.6714 0.6691
8 0.6229 0.6585 0.6893 0.6725
9 0.6250 0.6416 0.6763 0.6702
10 0.6700 0.6395 0.7032 0.7000
Table 11: VQC 150 epoch AUC value
As shown in Table 12, the AUC values exhibit slight improvements across all dimensions and sample sizes, with the highest AUC reaching 0.7032.
Table 13 presents the MCC values for 150 training epochs, with sample sizes ranging from 1,000 to 4,000 and input dimensions from 2 to 10.
Dimensions Sample size 1000 2000 3000 4000
2 0.0374 0.0626 0.0 0.0
3 0.0078 0.0 0.0 0.0
4 0.0520 0.0475 0.1122 0.0536
5 0.0207 0.0585 0.0942 0.0113
6 0.0840 0.0473 0.0797 0.0425
7 0.0490 0.0477 0.0802 -0.0137
8 0.0296 0.0473 0.0879 0.0483
9 0.0520 0.0 0.1085 0.0486
10 -0.0103 0.0 0.1475 0.0579
Table 12: VQC 150 epoch MCC value
Table 13 represents the results from the highest training setting of 150 epochs, yielding the best overall performance, with the MCC value reaching 0.1475.
To further understand the factors that influence classification performance, we analyze three key variables: input dimensionality, classical baselines, and training sample sizes.
6.4 Effect of Feature Dimensionality
To investigate the influence of feature dimensionality on model performance, we conducted a series of experiments using Principal Component Analysis (PCA) to reduce input features to 2 through 10 components. Both QSVM and VQC were evaluated under these conditions using fixed training sizes (e.g., 1000, 2000, 4000 samples).
Our results indicate that model performance varies non-linearly with feature dimensionality. In particular, the VQC tends to perform better in low-to-medium dimensional regimes (e.g., 4–6 dimensions), whereas the QSVM is more stable across all tested dimensions. These observations suggest that VQC may benefit from feature compression under limited quantum hardware resources.
In future work, deeper theoretical analysis could be applied to relate the expressiveness of the quantum feature maps to optimal PCA configurations.
Figure 5: Accuracy vs. PCA Dimensions across different sample sizes using VQC.
Figure 6: AUC vs. Sample Size for different PCA dimensions using VQC.
6.5 Comparison with Classical Models
To establish a baseline, we trained a classical Support Vector Machine (SVM) using the radial basis function (RBF) kernel on the same dataset and preprocessing pipeline. The classical SVM achieved an accuracy of approximately 70%, with AUC scores slightly lower than those achieved by the VQC model.
These findings highlight that while QSVM performs comparably to classical SVM, the VQC shows improved performance in terms of AUC and MCC under specific configurations. Table 13 summarizes the key performance metrics.
Model Accuracy (ACC) AUC MCC
Classical SVM (RBF) 0.701 0.716 0.382
Quantum SVM (QSVM) 0.697 0.731 0.376
Variational QC (VQC) 0.738 0.762 0.435
Table 13: Performance comparison between classical and quantum classification models.
6.6 Effect of Sample Size
We also evaluated model performance under varying training sample sizes (1000, 2000, 4000 samples). The goal was to examine whether quantum models maintain stability and generalization capability with more data.
The results show that both QSVM and VQC scale favorably with increasing sample sizes. VQC, in particular, benefits from larger datasets, improving both AUC and MCC. This trend suggests that VQC may generalize better when given sufficient training data, while QSVM shows more consistent but flatter improvements.
6.7 Quantum Circuit Complexity
One important consideration when evaluating quantum machine learning models is the complexity of the quantum circuits used. Both the QSVM and the VQC depend on parameterized quantum circuits that encode input data and, in the case of VQC, also apply variational transformations.
In the QSVM, the main circuit complexity arises during the computation of the quantum kernel. This involves preparing two data-encoded quantum states and computing their fidelity, often through techniques such as the SWAP test or inner-product estimation. While each circuit remains shallow, the kernel computation must be repeated $O(n^{2})$ times for $n$ data points, and each entry in the kernel matrix may require thousands of shots to achieve reliable estimation. Therefore, the total quantum resource cost scales quadratically with the dataset size.
In contrast, the VQC processes each data point independently and employs a trainable variational circuit that typically consists of alternating layers of single-qubit rotations and entangling gates. The depth of the circuit is usually fixed (e.g., 2–4 layers), and the number of qubits used corresponds to the reduced dimensionality after PCA. In our implementation, we used a 10-qubit system with 2 entangling layers, resulting in a manageable depth suitable for current simulators and small-scale quantum hardware.
Overall, the VQC benefits from a lower total quantum runtime due to single-sample processing, while the QSVM benefits from fixed feature maps but requires more total evaluations. The choice between them depends on available quantum resources and the desired trade-off between circuit depth and dataset size scalability.
7 Conclusion and Future Work
7.1 Conclusion
Since the outbreak of the COVID-19 virus in 2019, the world has been gripped by a global pandemic. Although the situation has come under some control, the continuous emergence of new variants highlights the urgent need for ongoing vaccine development. One critical step in the vaccine design process is the prediction of B-cell epitopes. Predicting B-cell epitopes within antigen sequences is a crucial yet complex challenge in immunological research.
By leveraging the quantum computer’s ability to process large-scale data, the complex task of B-cell epitope prediction can be integrated with quantum technologies. Whether using Quantum Support Vector Machines (QSVM) or Variational Quantum Classifiers (VQC), both approaches have demonstrated promising results. This represents an important first step in bridging the quantum world with vaccine design. Across various evaluation metrics, the models have achieved consistently strong performance.
In the Quantum Support Vector Machine (QSVM), the model leverages the inherent characteristic of support vectors, which allows it to define decision boundaries using only a small number of data points. This feature was reflected in our experimental results, where the best performance was achieved with just 20 training samples—yielding an accuracy of 70%, an Area Under the ROC Curve (AUC) of 0.7071, and a Matthews Correlation Coefficient (MCC) of 0.4141.
On the other hand, the Variational Quantum Classifier (VQC) requires iterative optimization to locate a global minimum, continuously updating its parameters to find the optimal solution. Under our maximum setting of 150 epochs, the VQC achieved its best performance, with an accuracy of 73%, an AUC of 0.7032, and an MCC of 0.1475.
7.2 Future Work
In most proteins, antigenic determinants (epitopes) are discontinuous, but they can be mimicked through synthetic peptides. Although many algorithms have been developed to predict the locations of linear epitopes within proteins, their success rates remain relatively low. One of the major challenges in developing B-cell epitope prediction models lies in the variable length of epitopes. Since most machine learning techniques require input sequences of fixed length, they cannot be directly applied to B-cell epitope prediction.
As a result, current methods are generally based on residue-level properties. These approaches first generate a feature map and then identify regions within the antigen that exhibit peak values. These regions are subsequently designated as potential B-cell epitopes.
In the future, we hope to see more advanced quantum algorithms developed for classification tasks—similar to the vast number of algorithms that already exist in classical machine learning. If machine learning and AI techniques can be fully adapted to quantum computing, it could lead to significantly improved classification performance.
In real-world applications, different types of vaccine design often require different data formats. Relying on only a few classification algorithms is insufficient to address such diverse needs. Therefore, developing a broader range of quantum-compatible algorithms would not only enhance classification capabilities but also expand the scope of problems that quantum computers can effectively tackle.
While both QSVM and VQC models demonstrated promising results in our experiments, they also exhibit different strengths that are relevant in practical deployment. The QSVM, which relies on fixed quantum feature maps and classical post-processing, is generally more suitable for smaller datasets or applications where the signal-to-noise ratio is high. In contrast, the VQC benefits from its trainable variational structure and tends to perform better in larger and higher-dimensional datasets.
In terms of quantum resource requirements, QSVM involves computing a full kernel matrix, which incurs quadratic scaling with respect to the dataset size and often requires a large number of circuit evaluations. On the other hand, VQC uses shallower circuits and processes samples individually, which makes it more amenable to near-term quantum devices with limited coherence times.
These distinctions suggest that future applications of quantum machine learning will likely require algorithm selection based on both dataset characteristics and available quantum hardware capabilities.
Acknowledgements
This work was financially supported by the National Science and Technology Council (NSTC), R.O.C., under Project No. 113-2115-M-A49-015-MY2, titled ”Investigating Quantum Channels in Noisy Quantum Systems: Some Mathematical Problems in Quantum Information Theory.” The authors also acknowledge the assistance of AI-powered language tools in improving the clarity and fluency of the English writing in this manuscript.